# Fast Radio Bursts as a probe of gravity on cosmological scales

### v1

@@ Appendix A Accuracy and details of the cosmopower emulator
An analysis of the test data reveals the emulators to be accurate within 5% (see Fig. A.1 for more information).
@@
for each mode, $k$ or $z$, depending on the function. Finally, by taking a histogram of the maximum and mean over all modes for each set of input parameters, a visual representation of how the neural network performs at worst and on average is created, shown in Fig. A.1. At worst, the emulator deviates $\leq 3\%$ and performs much better most of the time. Therefore, we exclude the emulator as a dominating systematic in this work.
Appendix B MCMC convergence test
The convergence tests are conducted by varying the hyperparameters of Nautilus-sampler: the maximum fraction of the evidence contained in the live set before the exploration phase ends, “f_live”, and the minimum demanded effective sample size, “n_eff”. In Fig B.1 we show the results for the Euclid-CS auto-correlations MCMCs with the fiducial cosmology given in Tab 3. Note that varying the hyperparameters has very little effect on the resulting contours, hence we use the Nautilus standard settings. Similar behaviour is obtained when testing the Euclid-CS×$\times$×FRB cross-correlation posterior for convergence.
Figure B.1: Euclid-CS auto-correlation posterior contours for different settings of Nautilus-sampler.

### v2

@@ Appendix B Accuracy and details of the cosmopower emulator
An analysis of the test data reveals the emulators to be accurate within 5% (see Fig. B.2 for more information).
Figure B.2: Histogram of the maximum and mean deviation of the emulated a) $P_{\mathrm{mm}}(k)$, b) $b_{\mathrm{e}}(k)$, c) $\eta(k)$, d) $\mu(k)$, e) $\chi(z)$ to the simulated test data.
@@
for each mode, $k$ or $z$, depending on the function. Finally, by taking a histogram of the maximum and mean over all modes for each set of input parameters, a visual representation of how the neural network performs at worst and on average is created, shown in Fig. B.2. At worst, the emulator deviates $\leq 3\%$ and performs much better most of the time. Therefore, we exclude the emulator as a dominating systematic in this work.
Appendix C MCMC convergence test
The convergence tests are conducted by varying the hyperparameters of Nautilus-sampler: the maximum fraction of the evidence contained in the live set before the exploration phase ends, “f_live”, and the minimum demanded effective sample size, “n_eff”. In Fig C.1 we show the results for the Euclid-CS auto-correlations MCMCs with the fiducial cosmology given in Tab 2. Note that varying the hyperparameters has very little effect on the resulting contours, hence we use the Nautilus standard settings. Similar behaviour is obtained when testing the Euclid-CS×$\times$×FRB cross-correlation posterior for convergence.
Figure C.1: Euclid-CS auto-correlation posterior contours for different settings of Nautilus-sampler.
Appendix D FRBs used for real life application
In Tab. 5 we summarise all the FRBs used in Section 5:
FRB $\mathrm{DM}_{\mathrm{obs}}$ z $\mathrm{DM}_{\mathrm{MW}}$ ra[rad] dec[rad] Source
20220207C 262.38 0.04304 79.3 -2.7807 1.272 Sherman et al. (2024)
20220307B 499.27 0.248123 135.7 -2.7333 1.26 Sherman et al. (2024)
20220310F 462.24 0.477958 45.4 -2.9848 1.2827 Sherman et al. (2024)
20220418A 623.25 0.622 37.6 -2.8867 1.2234 Sherman et al. (2024)
20220506D 396.97 0.30039 89.1 -2.7715 1.2711 Sherman et al. (2024)
20220825A 651.24 0.241397 79.7 -2.7786 1.2668 Sherman et al. (2024)
20220920A 314.99 0.158239 40.3 -2.862 1.2378 Sherman et al. (2024)
20221012A 441.08 0.284669 54.4 -2.8149 1.2309 Sherman et al. (2024)
20220914A 631.29 0.1139 47.55 1.7812 1.28 Sherman et al. (2024)
20220509G 269.53 0.0894 46.55 1.7919 1.226 Sherman et al. (2024)
20190608A 339.5 0.1178 204.75 2.6882 -0.1378 CHIME/FRB Collaboration et al. (2021)
20200430A 380.1 0.1608 112.21 0.8675 0.216 Kumar et al. (2021)
20150517A 560.0 0.19273 37.53 -1.6929 0.5783 Spitler et al. (2016)
20191001A 506.92 0.234 24.87 2.502 -0.9555 Shannon et al. (2019)
20190714A 504.7 0.2365 154.47 0.0695 -0.2273 Bhandari et al. (2019)
20190102C 363.6 0.291 17.67 2.4856 -1.3871 Macquart et al. (2020)
20180924B 362.4 0.3212 34.43 2.55 -0.7138 Bannister et al. (2019)
20200906A 577.8 0.3688 91.27 -2.2079 -0.2458 Bhandari et al. (2022)
20190611B 321.4 0.378 17.76 2.4565 -1.3857 Macquart et al. (2020)
20181112A 589.27 0.4755 25.76 2.5717 -0.9245 Prochaska et al. (2019)
20190523A 760.8 0.66 18.76 0.4724 1.2648 Ravi et al. (2019)
Table 5: FRBs used for the analysis presented in Section 5.